lexxii:

Show that this function is one-to-one algebraically. f(x)=(x−2)^3+8

Hero:

@lexxii, First replace f(x) with y

Hero:

Can you do that?

lexxii:

y=(x-2)^3+8

Hero:

Now swap the x and the y.

lexxii:

x=(y-2)^3+8

Hero:

Subtract 8 from both sides

lexxii:

x-8=(y-2)^3+8-8 x-8=(y-2)^3 ?

Hero:

Very good. Now apply the cube root to both sides.

lexxii:

Would the cube root go to the x or?

Hero:

It goes over the entire expression on either side.

lexxii:

so (x-8)^3=(y-2)^3 ?

Hero:

Well, you "cubed" one side and left the other sided "cubed". but you did not apply the "cube root" to both sides.

Hero:

I recommend you either use \(\LaTeX\) or the drawing feature to write your steps.

lexxii:

 ?

Hero:

Almost. We want to eliminate the cubed term on the RHS.

Hero:

To do this we apply the cube root. So it should be \(\sqrt[3]{x - 8} = \sqrt[3]{(y - 2)^3}\)

Hero:

When done in this manner, the cube and cube root cancel on the RHS

lexxii:

oh okay

Hero:

Actually we're finding the inverse of the function using this method. It seems that to determine whether the function is one-to one, we have to first re-write the function in terms of f(a) and f(b) then set the expressions equal to each other.

Hero:

So what we should do is: \(f(a) = (a - 2)^3 + 8\) \(f(b) = (b - 2)^3 + 8\) Then set \(f(a) = f(b)\) to get \((a - 2)^3 + 8 = (b - 2)^3 + 8\) And then keep applying inverse operations until we get \(a = b\)

lexxii:

Okay that makes sense I can finish it on my own from there thank you so much!

Hero:

You're welcome